Solve by substitution. 4x – 3y = 5 and x + 2y = 4
Answer by Boreal(15235)

4x – 3y = 5 and x + 2y = 4; use the second, because x stands alone. x=-2y+4
4(-2y+4)-3y=5
-8y+16-3y=5
-11y=-11
y=1
x=2
(2, 1)
Substitute into both equations to show that it checks.